# Zhixuan Chu1∗♠, Yan Wang1∗, Longfei Li1, Zhibo Wang2, Zhan Qin2, Kui Ren2

### v1

@@ Abstract
Large Language Models (LLMs) have shown impressive performance in natural language tasks, but their outputs can exhibit undesirable attributes or biases. Existing methods for steering LLMs towards desired attributes often assume unbiased representations and rely solely on steering prompts. However, the representations learned from pre-training can introduce semantic biases that influence the steering process, leading to suboptimal results. We propose LLMGuardaril, a novel framework that incorporates causal analysis and adversarial learning to obtain unbiased steering representations in LLMs. LLMGuardaril systematically identifies and blocks the confounding effects of biases, enabling the extraction of unbiased steering representations. Additionally, it includes an explainable component that provides insights into the alignment between the generated output and the desired direction. Experiments demonstrate LLMGuardaril’s effectiveness in steering LLMs towards desired attributes while mitigating biases. Our work contributes to the development of safe and reliable LLMs that align with desired attributes. We discuss the limitations and future research directions, highlighting the need for ongoing research to address the ethical implications of large language models.
**footnotetext: These authors contributed equally to this work.$\spadesuit$\$$\spadesuit$\$footnotetext: Corresponding author.
1 Introduction
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities in natural language understanding and generation, enabling a wide range of applications such as dialogue systems, content creation, time series forecasting, recommendation systems, professional agents [1; 2; 3; 4; 5]. However, the training of these models on massive web-scraped corpora has led to the manifestation of undesirable behaviors, such as generating offensive, toxic, or false outputs, facilitating the spread of disinformation campaigns. Addressing these issues is crucial, as content safety, fairness, toxicity, harmfulness, and factuality demand rigorous consideration, especially in an era where language models are increasingly deployed in high-stakes applications and user-facing systems. To address these challenges, recent research efforts have focused on developing methods to steer the output of LLMs towards desired attributes or concepts.
@@
Several approaches have been proposed to control and steer the generation process, aiming to mitigate these harmful behaviors and improve the overall quality and safety of generated content [6]. Fine-tuning on carefully curated datasets has been a common technique to enhance content safety and regulate outputs. By training language models on datasets that have been filtered and cleaned to remove offensive, biased, or misleading content, researchers aim to reduce the likelihood of the model generating such harmful outputs. In addition to fine-tuning, other techniques have been explored to further improve the generation process. Reinforcement learning from human feedback (RLHF) [7] is one such approach, where the language model is trained using feedback from human annotators who evaluate the quality and safety of generated outputs. Another promising approach is reinforcement learning from AI feedback (RLAIF) [8], which extends the concept of RLHF by using an AI system to provide feedback instead of human annotators. However, they all require huge annotation and computation resources. Furthermore, the training process often involves a human or AI annotator providing feedback and guidance to the language model. This raises the possibility that some form of deception or manipulation could be introduced during the training process, either intentionally or unintentionally [9; 10; 11]. In addition, these methods often lack explainability and interpretability, leading to inconsistent performance and limited generalizability [12; 13; 14; 15].
@@
Recent studies have shed light on the rich semantic information encoded within LLM representations, including specific information [16], concepts [17; 18], or attributes [19; 20]. Building upon these findings, the activation engineering has emerged, which involves creating vectors of activations that, when added to the forward passes of a frozen LLM, can induce desired changes in the output text. This approach holds promise for controlling the generation process in a more interpretable and fine-grained manner. Existing approaches for steering LLMs typically involve constructing positive and negative prompts that represent the desired and undesired attributes, respectively. By comparing the activations of the model for these prompts, a steering vector is obtained, which is then used to guide the model’s output during inference. While these methods have shown promising results, they often rely on the assumption that the constructed prompts are sufficient to capture the desired steering direction and that the model’s representations are unbiased. However, a closer examination reveals that the representations learned by LLMs during pre-training can introduce biases that influence the steering process. As shown in Figure 1, the semantic context of the prompts used for steering can implicitly encode biases, even without explicit steering prompts. Consequently, the obtained steering vectors may be confounded by these inherent biases, leading to suboptimal or unintended steering results. Therefore, simply structuring a controlled trial superficially is not enough to get the unbiased steering representation.
@@
Figure 1: Proportion of representations of semantic prompts that implicitly encode positive or negative directions for semantically neutral prompts without explicit steering prompt, across different language models. The varying proportions of positive and negative directions learned by the probing classifier, even in the absence of steering prompts, demonstrate the presence of inherent biases in the representations of semantic prompts due to differences in pre-training data. This observation supports the existence of a direct edge from the semantic direction representation $R_{cd}$ of the semantic prompt to the direction representation $R_{+}/R_{-}$, as discussed in the causal analysis section.
@@
To address these limitations, we propose a novel framework called LLMGuardaril that incorporates causal analysis to obtain unbiased steering representations for LLMs. Our framework aims to disentangle the influence of biases from the steering process by employing adversarial learning techniques. By systematically identifying and blocking the confounding effects of biases, LLMGuardaril enables the extraction of steering representations that accurately capture the desired attributes or concepts. The key contributions of this work are as follows:
@@
  • A novel framework for obtaining unbiased steering representations in LLMs. Our framework employs adversarial learning techniques to disentangle the influence of semantic biases from the steering process, enabling the extraction of accurate steering representations.
@@
  • Comprehensive experiments and analysis demonstrating the effectiveness of LLMGuardaril in steering LLMs towards desired attributes while mitigating the influence of biases. We evaluate our framework on various benchmark datasets and compare its performance with existing methods.
2 Background
2.1 Representations in Large Language Models
@@
Large language models (LLMs) have demonstrated remarkable capabilities in natural language processing tasks, generating human-like text and exhibiting a broad understanding of language. Central to their success is the rich set of representations they learn during training, which encode various concepts, attributes, and semantic information. Extensive research has been dedicated to understanding the representations learned by large language models (LLMs) and how they encode various concepts and attributes [18; 21; 22; 23]. LLM representations refer to the patterns of activations across the model’s parameters that correspond to specific semantic concepts, properties, or features. These representations act as the model’s internal codification of the information learned from the training data. Recent studies have provided compelling evidence that LLM representations contain rich semantic information, including abstract concepts like space and time [16], as well as more granular attributes related to truthfulness, toxicity, bias, and harmfulness of the generated text [24; 20; 25; 26]. In addition, linear classifier probes, which are trained to predict input properties from intermediate layers of a network, have successfully identified representations of concepts [27; 28]. Latent space analysis has enabled researchers to locate or edit factual associations within LLMs [29; 30; 31]. The ability to locate and manipulate these representations opens up avenues for controlling and steering the model’s outputs, enabling the development of principled methods for mitigating harmful behaviors and enhancing the interpretability and trustworthiness of large language models.
2.2 LLM Activation Engineering
@@
Activation engineering is a set of techniques that modify the internal activations of a pre-trained language model during inference to steer its output in a desired direction. The main idea is to identify and manipulate specific activations or attention heads associated with particular attributes or behaviors to control the model’s generation process. Early approaches, like the Plug-and-Play Language Model [32], use a separate classifier to detect target attributes in the generated text and perturb the language model’s activations accordingly, encouraging it to generate text that aligns with the desired attribute. Recent work focuses on extracting latent steering vectors from a frozen language model, which can be added to the activations during inference to steer the model’s completions toward specific goals, such as achieving high BLEU scores [33] or generating truthful statements [20]. Instead of requiring additional optimization or labeled data, Activation Addition [34] takes activation differences resulting from pairs of prompts. To avoid the identification of an opposite behaviour, [35] takes the average of activations associated with a target dataset and then subtracts the mean of all training activations, resulting in effective steering vectors. The growing interest in activation engineering stems from the desire to control and improve the performance of large language models on specific tasks or domains by identifying and manipulating the relevant activations.
2.3 Causal Inference
Figure 2: The causal graph and backdoor adjustment.
@@ 2.3 Causal Inference
Causal inference [36; 37; 38; 39; 40] has been an attractive research topic for a long time as it provides an effective way to uncover causal relationships in real-world problems. Nowadays, causal inference has shown potential in enhancing LLMs from a causal view in improving the LLMs’ reasoning capacity [1; 41; 42; 43; 44], addressing fairness issues in LLMs [45; 46; 47; 44], and safety [48; 49; 50], complementing LLMs with explanations [51; 13; 52; 53], and handling multimodality [54; 55]. In causal inference, if a variable is the common cause of two variables, it is called the confounder. The confounder will induce a spurious correlation between these two variables to disturb the recognition of the causal effect between them [56]. As shown in Figure 2(a), $X\rightarrow Y$ denotes that $X$ is the cause of $Y$. $C$ is the cause of both $X$ and $Y$. Thus, it is a confounder that will induce a spurious correlation between $X$ and $Y$ to disturb the recognition of the causal effect between them. In particular, such spurious correlation is brought by the backdoor path created by the confounder. Formally, a backdoor path between $X$ and $Y$ is defined as any path from $X$ to $Y$ that starts with an arrow pointing into $X$. For example, the path $X\leftarrow C\rightarrow Y$ is a backdoor path. If we want to deconfound two variables $X$ and $Y$ to calculate the true causal effect, we should block every backdoor path between them [57]. For example, in Figure 2(b), we should block $X\leftarrow C\rightarrow Y$ to get the causal effect between $X$ and $Y$.
3 Causal Analysis
3.1 The Hypothetical Situation
@@
Recent activation engineering-based work [26; 18; 34; 35; 20] utilize the steering vectors to control the direction of LLM output by constructing the randomized controlled trials. For example, some work [26; 18; 34] construct a pair of natural-language prompts ($p_{+}$, $p_{-}$), where $p_{+}$ represents the attribute we wish output text to emphasize and $p_{-}$ represents its opposite. $R_{+}/R_{-}$ is the representation for the prompt $p_{+}/p_{-}$. The difference $\Delta R$ is a new steering vector that (intuitively) captures the difference between a prompt with the attribute and without it. To obtain a steering vector, they perform a forward pass on each prompt, record the activations at the given location in each pass, take the difference, and then finally rescale this difference in activations by an “injection coefficient” $\beta$. To steer, they add the resulting steering representation to the original representations and allow the forward pass to continue, and obtain the steered output.
@@
To systemically analyze this series of methods, we give the formal definitions. Now, let’s further break down and analysis of the constructed prompt pairs. In fact, the input prompt is comprised of the semantic prompt $C$ and the steering prompt $S$. For example, in the [34], “I love talking about weddings” is the positive prompt and “I hate talking about weddings” is the negative prompt. In these two examples, “love” and “hate” are the steering prompts, which control the directions of output. The “I … talking about weddings” is the same for this pair of prompts, which only contains the semantic information. Therefore, they assume this is a randomized controlled trial, where there are two groups, i.e., treatment group “love” and control group “hate”. As shown in Figure 3 (a), the semantic prompt $C$ is a confounder, that can affect the steering prompt $S$ and output $Y$. In this hypothetical situation, they construct the randomized controlled trial by creating the pairs, i.e., positive steering prompt plus the same semantic prompt and negative steering prompt plus the same semantic prompt. The only difference is the steering prompt since the semantic prompts are the totally same. Based on this assumption, the edge from the semantic prompt to the steering prompt is blocked. Therefore, they can get the difference in activations, which can intervene in the generated output. The strength of this steering vector is called the treatment effect in causal inference. On the face of it, this is a perfect randomized controlled trial, where only the steering prompt can affect the direction representation $R_{+}/R_{-}$. The combination of direction representation $R_{+}/R_{-}$ and the semantic prompt together have an influence on the generated output $Y$. In this case, the steering vector can be obtained by the operation between positive direction representation $R_{+}$ and negative direction representation $R_{-}$.
Figure 3: The causal analysis of our proposed LLMGuardaril.
3.2 The Real Situation

### v2

@@ Abstract.
Large Language Models (LLMs) have shown impressive performance in natural language tasks, but their outputs can exhibit undesirable attributes or biases. Existing methods for steering LLMs toward desired attributes often assume unbiased representations and rely solely on steering prompts. However, the representations learned from pre-training can introduce semantic biases that influence the steering process, leading to suboptimal results. We propose LLMGuardrail, a novel framework that incorporates causal analysis and adversarial learning to obtain unbiased steering representations in LLMs. LLMGuardrail systematically identifies and blocks the confounding effects of biases, enabling the extraction of unbiased steering representations. Additionally, it includes an explainable component that provides insights into the alignment between the generated output and the desired direction. Experiments demonstrate LLMGuardrail’s effectiveness in steering LLMs toward desired attributes while mitigating biases. Our work contributes to the development of safe and reliable LLMs that align with desired attributes.
Large Language Model, Causal Inference, Explanation, Trustworthiness, Safety
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security; October 14–18, 2024; Salt Lake City, UT, USAbooktitle: Proceedings of the 2024 ACM SIGSAC Conference on Computer and Communications Security (CCS ’24), October 14–18, 2024, Salt Lake City, UT, USAdoi: 10.1145/3658644.3690217isbn: 979-8-4007-0636-3/24/10ccs: Computing methodologies Natural language generationccs: Computing methodologies Causal reasoning and diagnostics
1. Introduction
@@
Large Language Models (LLMs) have demonstrated remarkable capabilities in natural language understanding and generation, enabling a wide range of applications such as dialogue systems, clinical report generation, professional agents, recommendation systems, and so on (Liu et al., 2024; Guan et al., 2024; Chu et al., 2024c, b; Wang et al., 2024b). However, the training of these models on massive web-scraped corpora has led to the manifestation of undesirable behaviors, such as generating offensive, toxic, or false outputs (Huang et al., 2023; Wen et al., 2023; Chu et al., 2024d). Addressing these issues is crucial, as content safety, fairness, toxicity, harmfulness, and factuality demand rigorous consideration, especially in an era where language models are increasingly deployed in high-stakes applications and user-facing systems.
@@
To address these challenges, recent research efforts have focused on developing methods to steer the output of LLMs toward desired attributes or concepts. Several approaches have been proposed to control and steer the generation process, aiming to mitigate these harmful behaviors and improve the overall quality and safety of generated content. Fine-tuning on carefully curated datasets has been a common technique to enhance content safety and regulate outputs. By training language models on datasets that have been filtered and cleaned to remove offensive, biased, or misleading content, researchers aim to reduce the likelihood of the model generating such harmful outputs. In addition to fine-tuning, other techniques have been explored to further improve the generation process. Reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) is one such approach, where the language model is trained using feedback from human annotators who evaluate the quality and safety of generated outputs. Another promising approach is reinforcement learning from AI feedback (RLAIF) (Lee et al., 2023a), which extends the concept of RLHF by using an AI system to provide feedback instead of human annotators. However, they all require huge annotation and computation resources. Furthermore, the training process often involves a human or AI annotator providing feedback and guidance to the language model. This raises the possibility that some form of deception or manipulation could be introduced during the training process, either intentionally or unintentionally (Tarsney, 2024; Tan et al., 2024; Qi et al., 2023). In addition, these methods often lack explainability and interpretability, leading to inconsistent performance and limited generalizability (Liao and Vaughan, 2023; Zhao et al., 2024; Hickling et al., 2023).
@@
Recent studies have shed light on the rich semantic information encoded within LLM representations, including specific information (Gurnee and Tegmark, 2023), concepts (Mikolov et al., 2013; Zou et al., 2023a), or attributes (Bolukbasi et al., 2016; Li et al., 2024). Building on these findings, activation engineering has emerged, which induces desired outputs in frozen LLMs by injecting crafted activation vectors during forward passes. This approach holds promise for controlling the generation process in a more interpretable and fine-grained manner. Existing approaches for steering LLMs typically involve constructing positive and negative prompts that represent the desired and undesired attributes, respectively. By comparing the activations of the model for these prompts, a steering vector is obtained, which is then used to guide the model’s output during inference. While these methods have shown promising results, they often rely on the assumption that the constructed prompts are sufficient to capture the desired steering direction and that the model’s representations are unbiased. However, a closer examination reveals that the representations learned by LLMs during pre-training can introduce biases that influence the steering process. As shown in Figure 1, the semantic context of the prompts used for steering can implicitly encode biases, even without explicit steering prompts. Consequently, the obtained steering vectors may be confounded by these inherent biases, leading to suboptimal or unintended steering results. This observation highlights a critical gap in current steering approaches: the need to account for and mitigate the influence of semantic biases on the steering process. Without addressing this issue, steering methods may inadvertently perpetuate or even amplify existing biases in the model’s outputs, undermining efforts to improve the safety and reliability of LLMs. Therefore, simply structuring a controlled trial superficially is not enough to get an unbiased steering representation.
@@
Figure 1. Proportion of representations of semantic prompts that implicitly encode positive or negative directions for semantically neutral prompts without explicit steering prompt, across different language models. The varying proportions of positive and negative directions learned by the probing classifier, even in the absence of steering prompts, demonstrate the presence of inherent biases in the representations of semantic prompts due to differences in pre-training data. This observation supports the existence of a direct edge from the semantic direction representation $R^{cd}$ of the semantic prompt to the direction representation $R_{+}/R_{-}$, as discussed in the causal analysis section.
@@
To address these limitations, we propose a novel framework called LLMGuardrail that incorporates causal analysis to obtain unbiased steering representations for LLMs. Our framework aims to disentangle the influence of biases from the steering process by employing debiasing techniques. By systematically identifying and blocking the confounding effects of biases, LLMGuardrail enables the extraction of steering representations that accurately capture the desired attributes. The key contributions of this work are as follows:
@@
  • A novel framework for obtaining unbiased steering representations in LLMs. Our framework employs adversarial learning techniques to disentangle the influence of semantic biases from the steering process.
  • An explainable component that provides insights into the alignment between the generated output and the desired direction, enhancing the interpretability of the steering process.
@@
  • Comprehensive experiments and analysis demonstrating the effectiveness of LLMGuardrail in steering LLMs toward desired attributes while mitigating the influence of biases. We evaluate our framework on various benchmark datasets and compare its performance with existing methods.
Figure 2. The causal analysis of our proposed LLMGuardrail. (a) The Hypothetical situation: The constructed pair of prompts can block the edge from the semantic prompt $C$ to the steering prompt $S$ and the direction representation $R^{+}/R^{-}$. (b) The Real situation: In addition to the edge from the semantic prompt $C$ to the steering prompt $S$, there is a direct edge from the semantic direction representation $R^{cd}$ of the semantic prompt to the direction representation $R^{+}/R^{-}$. (c) Ours: We need to block the edge from the semantic direction representation $R^{cd}$ of the semantic prompt to the direction representation $R^{+}/R^{-}$. The direction representation $R^{+}/R^{-}$ is only influenced by the steering prompt $S^{+}/S^{-}$, enabling us to obtain an unbiased steering representation $\Delta R$ through steering engineering.
2. Background
2.1. Representations in Large Language Models
@@
Large language models (LLMs) have demonstrated remarkable capabilities in natural language processing tasks, generating human-like text and exhibiting a broad understanding of language. Central to their success is the rich set of representations they learn during training, which encode various concepts, attributes, and semantic information. Extensive research has been dedicated to understanding the representations learned by large language models (LLMs) and how they encode various concepts and attributes (Zou et al., 2023a). LLM representations refer to the patterns of activations across the model’s parameters that correspond to specific semantic concepts, properties, or features. These representations act as the model’s internal codification of the information learned from the training data. Recent studies have provided compelling evidence that LLM representations contain rich semantic information, including abstract concepts like space and time (Gurnee and Tegmark, 2023), as well as more granular attributes related to truthfulness, toxicity, bias, and harmfulness of the generated text (Chen et al., 2024b; Li et al., 2024; Azaria and Mitchell, 2023; Burns et al., 2022). In addition, linear classifier probes, which are trained to predict input properties from intermediate layers of a network, have successfully identified representations of concepts (Alain and Bengio, 2016; Belinkov, 2022). Latent space analysis has enabled researchers to locate or edit factual associations within LLMs (Hernandez et al., 2023; Meng et al., 2022; Zhong et al., 2023). The ability to locate and manipulate these representations opens up avenues for controlling and steering the model’s outputs, enabling the development of principled methods for mitigating harmful behaviors and enhancing the interpretability and trustworthiness of large language models.
2.2. LLM Activation Engineering
@@
Activation engineering is a set of techniques that modify the internal activations of a pre-trained language model during inference to steer its output in a desired direction. The main idea is to identify and manipulate specific activations or attention heads associated with particular attributes or behaviors to control the model’s generation process. Early approaches, like the Plug-and-Play Language Model (Dathathri et al., 2019), use a separate classifier to detect target attributes in the generated text and perturb the language model’s activations accordingly, encouraging it to generate text that aligns with the desired attribute. Recent work focuses on extracting latent steering vectors from a frozen language model, which can be added to the activations during inference to steer the model’s completions toward specific goals, such as achieving high BLEU scores (Subramani et al., 2022) or generating truthful statements (Li et al., 2024). Instead of requiring additional optimization or labeled data, Activation Addition (Turner et al., 2023) takes activation differences resulting from pairs of prompts. To avoid the identification of an opposite behavior, (Jorgensen et al., 2023) takes the average of activations associated with a target dataset and then subtracts the mean of all training activations, resulting in effective steering vectors. The growing interest in activation engineering stems from the desire to control and improve the performance of large language models on specific tasks or domains by identifying and manipulating the relevant activations.
2.3. Causal Inference
Figure 3. The causal graph and backdoor adjustment.
@@ 2.3. Causal Inference
Causal inference (Pearl, 2000; Rubin, 2005; Yao et al., 2021; Chu and Li, 2023; Cui et al., 2020) has been an attractive research topic for a long time as it provides an effective way to uncover causal relationships in real-world problems. Nowadays, causal inference has shown potential in enhancing LLMs from a causal view in improving the LLMs’ reasoning capacity (Chu et al., 2023; Gao et al., 2023; Zhang et al., 2022; Chen et al., 2024a; Liu et al., 2022), addressing fairness issues in LLMs (Stanovsky et al., 2019; Ding et al., 2022; Meade et al., 2021), and safety (Bao et al., 2021; Cao et al., 2022; Liu et al., 2023a), complementing LLMs with explanations (Liu et al., 2023b; Zhao et al., 2024; Belrose et al., 2023; Wan et al., 2024), and handling multimodality (Ko et al., 2023; Pawlowski et al., 2023). In causal inference, if a variable is the common cause of two variables, it is called the confounder. The confounder will induce a spurious correlation between these two variables to disturb the recognition of the causal effect between them. As shown in Figure 3(a), $X\rightarrow Y$ denotes that $X$ is the cause of $Y$. $C$ is the cause of both $X$ and $Y$. Thus, it is a confounder that will induce a spurious correlation between $X$ and $Y$ to disturb the recognition of the causal effect between them. In particular, such spurious correlation is brought by the backdoor path created by the confounder. Formally, a backdoor path between $X$ and $Y$ is defined as any path from $X$ to $Y$ that starts with an arrow pointing into $X$. For example, the path $X\leftarrow C\rightarrow Y$ is a backdoor path. If we want to deconfound two variables $X$ and $Y$ to calculate the true causal effect, we should block every backdoor path between them (Pearl and Mackenzie, 2018). For example, in Figure 3(b), we should block $X\leftarrow C\rightarrow Y$ to get the causal effect between $X$ and $Y$.
3. Causal Analysis
3.1. The Hypothetical Situation
@@
Recent activation engineering-based work (Burns et al., 2022; Zou et al., 2023a; Turner et al., 2023; Jorgensen et al., 2023; Li et al., 2024) utilize the steering vectors to control the direction of LLM output by constructing the randomized controlled trials. For example, some work (Burns et al., 2022; Zou et al., 2023a; Turner et al., 2023) construct a pair of natural-language prompts ($p_{+}$, $p_{-}$), where $p_{+}$ represents the attribute we wish output text to emphasize and $p_{-}$ represents its opposite. $R_{+}/R_{-}$ is the representation for the prompt $p_{+}/p_{-}$. The difference $\Delta R$ is a new steering vector that (intuitively) captures the difference between a prompt with the attribute and without it. To obtain a steering vector, they perform a forward pass on each prompt, record the activations at the given location in each pass, take the difference, and then finally rescale this difference in activations by an “injection coefficient” $\beta$. To steer, they add the resulting steering representation to the original representations and allow the forward pass to continue, and obtain the steered output.
@@
To systemically analyze this series of methods, we give the formal definitions. Now, let’s further break down and analysis of the constructed prompt pairs. In fact, the input prompt is comprised of the semantic prompt $C$ and the steering prompt $S$. For example, in the (Turner et al., 2023), “I love talking about weddings” is the positive prompt and “I hate talking about weddings” is the negative prompt. In these two examples, “love” and “hate” are the steering prompts, which control the directions of output. The “I … talking about weddings” is the same for this pair of prompts, which only contains the semantic information. Therefore, they assume this is a randomized controlled trial, where there are two groups, i.e., treatment group “love” and control group “hate”. As shown in Figure 2 (a), the semantic prompt $C$ is a confounder, that can affect the steering prompt $S$ and output $Y$. In this hypothetical situation, they construct the randomized controlled trial by creating the pairs, i.e., positive steering prompt plus the same semantic prompt and negative steering prompt plus the same semantic prompt. The only difference is the steering prompt since the semantic prompts are the totally same. Based on this assumption, the edge from the semantic prompt to the steering prompt is blocked. Therefore, they can get the difference in activations, which can intervene in the generated output. The strength of this steering vector is called the treatment effect in causal inference. On the face of it, this is a perfect randomized controlled trial, where only the steering prompt can affect the direction representation $R_{+}/R_{-}$. The combination of direction representation $R_{+}/R_{-}$ and the semantic prompt together have an influence on the generated output $Y$. In this case, the steering vector can be obtained by the operation between positive direction representation $R_{+}$ and negative direction representation $R_{-}$.
3.2. The Real Situation